OpenStudy (4meisu):

e^2x (lnx) - e^2x = 0 Solve.

sam (.sam.):

Is it \[e ^{2x}(lnx)-e^{2x}=0\] OR \[e ^{2}x(lnx)-e^{2}x=0\]

sam (.sam.):

e^(2x)(lnx)-e^(2x)=0 lnxe^(2x)-e^(2x)=0 lnxe^(2x)=e^(2x) lnx=1 e^(lnx)=e x=e

OpenStudy (4meisu):

It's the first one where the 2x is the exponent :)

OpenStudy (bahrom7893):

i see :)

sam (.sam.):

Yeah that's the solution

OpenStudy (4meisu):

Thank you!